OpenStudy (anonymous):

7(x -2) + 3(x + 2) = 5(x - 6) I need to really understand this..

OpenStudy (anikate):

i will help u

OpenStudy (anikate):

ready

OpenStudy (anonymous):

Sure thing :)

OpenStudy (anonymous):

First, distribute the numbers across your parentheses that they're attached to. This is as easy as multiplying them.

OpenStudy (anonymous):

7(x-2) will turn into 7x - 14 7 times x is 7x and 7 times negative 2 is negative 14

OpenStudy (anikate):

wait

OpenStudy (anikate):

first simplify the question

OpenStudy (anikate):

and write

OpenStudy (anonymous):

3(x+2) equals 3x + 6 and 5(x-6) is 5x - 30

OpenStudy (anikate):

together it looks like this

OpenStudy (anonymous):

So....... we have: 7x - 14 + 3x + 6 = 5x - 30

OpenStudy (anikate):

7x-14+3x+6=5x-30

OpenStudy (anikate):

now put the x's on the left side

OpenStudy (anonymous):

Here's where it gets a little tricky. Begin by adding numbers that don't have an x *If they're on the same side of the = sign*

OpenStudy (anikate):

and the normal numbers on the right

OpenStudy (anikate):

so it looks like this

OpenStudy (anikate):

7x+3x-5x=14-6-30

OpenStudy (anonymous):

So you have 7x + 3x - 8 = 5x - 30

OpenStudy (anikate):

now simplify it

OpenStudy (anonymous):

Now, do the same, but for numbers that have X values 10x - 8 = 5x -30

OpenStudy (anikate):

10x-5x=-30+8

OpenStudy (anikate):

what is is?

OpenStudy (anonymous):

Now, choose which side you want X values to be, and which side you want non-X values to be. I like putting X on the left. So... add 8 to both sides. This will 'kill' the 8 on the left and make the -30 turn into -22

OpenStudy (anonymous):

and you can now subtract 5x from both sides, making the 5x 'die' off, and turning your 10x into just 5x (10x-5x = 5x)

OpenStudy (anikate):

it would be 5x=-22

OpenStudy (anikate):

so the answer is x=-22/5

OpenStudy (anonymous):

then you are left with -22/5 = x And there you are :D

OpenStudy (anikate):

i was helping u

OpenStudy (anikate):

not the other way around

OpenStudy (anonymous):

Are you talking to me, Anikate?

OpenStudy (anikate):

understand the question now

OpenStudy (anikate):

yes

OpenStudy (anonymous):

Lol. I was helping ja93de on my own.

OpenStudy (anikate):

o

OpenStudy (anonymous):

So, do you understand, ja93de? Both our methods work, we just came to the final answer in a different order. ^_^

OpenStudy (anikate):

k well

OpenStudy (anonymous):

Medal please. :P